OpenStudy (anonymous):

the light from a 20 ft street lamp casts a shadow infront of a 5 3/4 foot tall person who is standing 24 feet from the lamppost. How long is the persons shadow

OpenStudy (zzr0ck3r):

Can you draw a picture including the measurements given?

OpenStudy (zzr0ck3r):

@blahblah212121 ?????

OpenStudy (anonymous):

@zzr0ck3r there are no pictures

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

right

OpenStudy (zzr0ck3r):

do you know about similar triangles?

OpenStudy (anonymous):

yes kinda

OpenStudy (zzr0ck3r):

note that \(\frac{20}{24+?}=\frac{5\frac{3}{4}}{?}\iff \frac{20}{24+x}=\frac{5\frac{3}{4}}{x}\). Can you solve for \(x\)

OpenStudy (anonymous):

yes thank you all i needed was the setup!